x2y''+xy'+4y=2xlnx


1
Expert's answer
2021-08-11T19:07:46-0400
"x^2y''+xy'+4y=2x\ln x"

Homogeneous equation


"x^2y''+xy'+4y=0"


Let "y=x^r"


"y'=rx^{r-1}"


"y''=r(r-1)x^{r-2}"


"r(r-1)+r+4=0"

"r=\pm2 i"

The general solution of the homogeneous equation is


"y_h=C_1\cos(2\ln x)+C_2\sin(2\ln x)"


"x^2y''+xy'+4y=2x\ln x"

Divide both sides by "x^2"


"y''+\dfrac{1}{x}y'+\dfrac{4}{x^2}y=\dfrac{2}{x}\ln x=>g(x)=\dfrac{2}{x}\ln x"

Wronskian


"W(y_1, y_2)=\begin{vmatrix}
 \cos(2\ln x)& \sin(2\ln x) \\
 - \dfrac{2}{x}\sin (2\ln x) & \dfrac{2}{x}\cos(2\ln x)
\end{vmatrix}=\dfrac{2}{x}"

"W_1=\begin{vmatrix}
 0& \sin(2\ln x) \\
 \dfrac{2}{x}\ln x & \dfrac{2}{x}\cos(2\ln x)
\end{vmatrix}=-\dfrac{2}{x}\ln x \sin(2\ln x)"


"W_2=\begin{vmatrix}
 \cos(2\ln x) & 0 \\
 - \dfrac{2}{x}\sin (2\ln x) & \dfrac{2}{x}\ln x
\end{vmatrix}=\dfrac{2}{x}\ln x \cos(2\ln x)"

"u_1=\int\dfrac{W_1}{W}dx=\int\dfrac{-\dfrac{2}{x}\ln x \sin(2\ln x)}{\dfrac{2}{x}}dx"

"=-\int \ln x \sin(2\ln x)dx"

"=-\dfrac{1}{25}x(5\ln x+3)\sin(2\ln x)+\dfrac{1}{25}(10\ln x-4)\cos(2\ln x)"


"u_2=\int\dfrac{W_2}{W}dx=\int\dfrac{\dfrac{2}{x}\ln x \cos(2\ln x)}{\dfrac{2}{x}}dx"

"=\int \ln x \cos(2\ln x)dx"

"=\dfrac{1}{25}x(10\ln x-4)\sin(2\ln x)+\dfrac{1}{25}(5\ln x+3)\cos(2\ln x)"

Particular solution of the non-homogeneous equation is


"y_p=u_1y_1+u_2y_2"

"y_p=-\dfrac{1}{25}x(5\ln x+3)\sin(2\ln x)\cos(2\ln x)"

"+\dfrac{1}{25}(10\ln x-4)\cos^2(2\ln x)"

"+\dfrac{1}{25}x(10\ln x-4)\sin^2(2\ln x)"

"+\dfrac{1}{25}(5\ln x+3)\sin(2\ln x)\cos(2\ln x)"

The general solution of the homogeneous equation is


"y=C_1\cos(2\ln x)+C_2\sin(2\ln x)"

"-\dfrac{1}{25}x(5\ln x+3)\sin(2\ln x)\cos(2\ln x)"

"+\dfrac{1}{25}(10\ln x-4)\cos^2(2\ln x)"

"+\dfrac{1}{25}x(10\ln x-4)\sin^2(2\ln x)"

"+\dfrac{1}{25}(5\ln x+3)\sin(2\ln x)\cos(2\ln x)"